OpenStudy (shiree19):

Please Help!! Fan and Medal! Problem - Solve using two different methods! Explain which method you found to be more efficient? A - 3x - 9y = 3 6x - 3y = -24 Thank you!

OpenStudy (shiree19):

@SolomonZelman @humphriescody

OpenStudy (anonymous):

I do not now sorey good luck thow

OpenStudy (shiree19):

Thanks anyway. :)

OpenStudy (shiree19):

@Loser66 @scburchell11

OpenStudy (shiree19):

Please help! Anyone?

OpenStudy (anonymous):

(a) elimination : 3x - 9y = 3 ---->(-2) 6x - 3y = -24 -------------- -6x + 18y = -6 (result of multiplying by -2) 6x - 3y = -24 -------------add 15y = - 30 y = - 2 sub in -2 for y 3x - 9y = 3 3x - 9(-2) = 3 3x + 18 = 3 3x = -18 + 3 3x = - 15 x = - 5 substitution 3x - 9y = 3 3x = 9y + 3 x = 9/3y + 3/3 x = 3y + 1 now sub 3y + 1 in for x 6x - 3y = -24 6(3y + 1) - 3y = -24 18y + 6 - 3y = -24 18y - 3y = -24 - 6 15y = -30 y = -2 now sub -2 in for y 3x - 9y = 3 3x - 9(-2) = 3 3x + 18 = 3 3x = 3 - 18 3x = -15 x = -5

OpenStudy (shiree19):

Ok but how did you get -2 for y? And did you use Substitution for both of them?

OpenStudy (anonymous):

explained in now sub 3y + 1 in for x

OpenStudy (shiree19):

So there would be 2 y's? That doesn't really make sense.

OpenStudy (shiree19):

Well if 3y + 1 = -2 = y is that right?

OpenStudy (anonymous):

yes

OpenStudy (shiree19):

Ok. So the answer for 3x - 9y = 3 is x = -5 & y = -2 and we used the Substitution method. Is that correct?

OpenStudy (anonymous):

yes

OpenStudy (shiree19):

So what method do we use to solve 6x - 3y = -24? As we can't use the same method.

OpenStudy (anonymous):

yes

OpenStudy (shiree19):

What method would be best? Maybe elimination?

OpenStudy (anonymous):

that is the answer for both

OpenStudy (shiree19):

Yes, but I have to use a different method for each problem.

OpenStudy (anonymous):

this one is the best I will try to finesh the uther one

OpenStudy (shiree19):

Using elimination - if y = -2 then to solve x 6x - 3 * -2 = -24 divide both sides by 6 x - 3 * -2 = -4 divide both sides by -2 x - 3 = -2 add 3 to both sides x = -5 So, x = -5 and y = -2 This is right, yes?

OpenStudy (anonymous):

that is it

OpenStudy (shiree19):

Ok. Thanks so much!

OpenStudy (anonymous):

ok you are welcome

OpenStudy (anonymous):

do you need more help

OpenStudy (shiree19):

Yes please. Sorry for the wait I was trying to figure it out on my own. Same rules apply. 7x - 3y = 20 5x + 3y = 16

OpenStudy (shiree19):

I think I figured it out.